OpenStudy (anonymous):

In slope-intercept form, write the equation of the line that is parallel to y=-2x-9 and passes through (-2,-4)

OpenStudy (cwrw238):

the line will have same slope as y = -2x - 9 use the general form y-y1 = m(x-x1) where m = -2, x1 = -2, y1 = -4 and simplify

OpenStudy (cwrw238):

m = slope = -2